How many grams of NaF would have to be added to 2.00 L of 0.100 M HF

to yield a solution with a pH = 4.00?


1
Expert's answer
2022-04-26T12:39:02-0400

pH = pKa + log ([F-]/[HF])


Ka(HF) = 6.8 * 10-4

pKa = - log Ka = - log 6.8 * 10-4 = 3.17


log ([F-]/[HF]) = pH - pKa

log ([F-]/0.1М) = 4 - 3.17 = 0.83

[F-]/0.1М = 100.83

[F-] = 0.68 M


m(NaF) = 0.68 M * 2 L * 41.99 g/mol = 57.10 g